OpenStudy (anonymous):

The leg of a right triangle is 4 units and the hypotenuse is 9 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (anonymous):

First of all, do you know what the Pythagorean theorem is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, well, the Pythagorean Theorem is a^2+b^2=c^2 (a and b are the legs of a triangle and c is the hypotenuse) Now, we know that the second leg has to be lower than 9 because the hypotenuse is always the longest on a right triangle. So, you would then plug the numbers into the equation. When you do that you'd have a^2+4^2=9^2 Do you know what 4^2 and 9^2 is?

OpenStudy (anonymous):

16 and 81

OpenStudy (anonymous):

@vangoo98

OpenStudy (anonymous):

Yes! 16 and 81. So, now you're equation would be a^2+16=81 Then you would subtract 16 from 81. What is 81-16?

OpenStudy (anonymous):

Your*

OpenStudy (anonymous):

65

OpenStudy (anonymous):

@vangoo98

OpenStudy (anonymous):

Yes, so, now you just take the square root of 65 and you have the other leg's length :) It will be a decimal because 65 is not a perfect square. Does that make sense?

OpenStudy (anonymous):

ohh yeah! thank you!

OpenStudy (anonymous):

You're welcome! :)